40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 293 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  2880   Wed May 5 01:19:05 2010 KojiUpdateIOOMC spot centering cont"d

Koji and Zach

We improved the beam axis rotaion on the MC. We still have 3mrad to be corrected.
So far we lost the MC Trans spot on CCD as the beam is now hitting the flange of the window. We need to move the steering mirror.

To do next:

- MC2 spot is too much off. Adjust it.

- Rotate axis for 3mrad more.

- MC2 spot is too much off. Adjust it.

- Adjust Vertical spot position as a final touch.


Monday

- Incident beam had 7mrad rotation.

- Tried to rotate in-vac steering mirror (IM1) in CCW

- After the long struggle the beam from PSL table started to hit north-east side of IM1 mount.

- Moved the IM1. All of the beam (input beam, MC Trans, MC Refl) got moved. Chaotic.

- Recovered TEM00 resonance. MC Trans CCD image missing. The beam axis rotation was 8.5mrad.
  Even worse. Disappointed.

Tuesday

- We made a strategic plan after some deliberation.

- We returned to the initial alignment of Saturday only for yaw.
  Not at once, such that we don't miss the resonance.

- Adjusted SM2Y and IM1Y to get reasonable resonance. Then adjusted MC2/3 to have TEM00 lock.

- Measured the spot positions. The axis rotation was 4.8mrad.

- Moved the spot on IM1 by 7mm by rotating SM2Y in CCW - ((A) in the figure)

- Compensated the misalignment by IM1Y CCW. ((B) in the figure)
  Used a large sensor card with puch holes to see the spot distribution between the MC1 and MC3.

- Fine alignment by MC2/MC3. Lock to TEM00. The beam axis rotation was 3mrad.The beam axis translation was 3mm.

- This 3mm can be Compensated by IM1Y. But this can easily let the resonance lost.
  Put the sensor card between MC1/MC3 and compensated the misalignment by MC3 and MC1.

Note: You match the returned spot from the MC2 to the incident beam by moving the spot deviation by MC3,
the spot returns to the good position on MC1. But the angle of the returned beam is totally bad.
This angle deviation can be adjusted by MC1.

Note2: This step should be done for max 2mm (2mrad) at once. As 2mrad deviation induces the spot move on the MC2 by an inch.

- After all, what we get is

MC1H = -0.15 mm
MC1V = -0.33 mm
MC3H = +0.97 mm
MC3V = -0.33 mm

This corresponds to the axis rotation of 3mrad and the beam axis translation of 0.8mm (to north).

Attachment 1: steering.png
steering.png
  2879   Tue May 4 18:40:27 2010 ranaHowToElectronicsMarconi phase noise measurement setup

To check the UGF, I increased the gain of the PLL by 10 and looked at how much the error point got suppressed. The green trace apparently has a UGF of ~50 Hz and so the BLUE nominal one has ~5 Hz.

The second attachment shows the noise now corrected for the loop gain. IF the two signal generators are equally noisy, then you can divide the purple spectrum by sqrt(2) to get the noise of a single source.

The .xml file is saved as /users/rana/dtt/MarconiPhaseNoise_100504.xml

Attachment 1: Untitled.png
Untitled.png
Attachment 2: ifrnoise.png
ifrnoise.png
  2878   Tue May 4 14:57:53 2010 josephbUpdateComputersOttavia has moved

Ottavia was moved this afternoon from the control room into the lab, adjacent to Mafalda in 1Y3 on the top shelf.  It has been connected to the camera hub, as well as the normal network.  Its cables are clearly labeled.  Note the camera hub cable should be plugged into the lower ethernet port. Brief tests indicate everything is connected and it can talk to the control room machines.

The space where Ottavia used to be is now temporarily available as a good place to setup a laptop, as there is keyboard, mouse, and an extra monitor available.  Hopefully this space may be filled in with a new workstation in the near future.

  2877   Tue May 4 13:14:43 2010 josephbUpdateCDSlsc.mdl and ifo.mdl to build (with caveats)

I got around to actually try building the LSC and IFO models on megatron.  Turns out "ifo" can't be used as a model name and breaks when trying to build it.  Has something to do with the find and replace routines I have a feeling (ifo is used for the C1, H1, etc type replacements throughout the code).  If you change the model name to something like ifa, it builds fine though.  This does mean we need a new name for the ifo model.

Also learned the model likes to have the cdsIPCx memory locations terminated on the inputs if its being used in a input role (I.e. its bringing the channel into the model).  However when the same part is being used in an output role (i.e. its transmitting from the model to some other model), if you terminate the output side, it gives errors when you try to make.

Its using the C1.ipc file (in /cvs/cds/caltech/chans/ipc/) just fine.  If you have missing memory locations in the C1.ipc file (i.e. you forgot to define something) it gives a readable error message at compile time, which is good.  The file seems to be being parsed properly, so the era of writing "0x20fc" for block names is officially over.

  2876   Tue May 4 06:32:58 2010 albertoConfigurationPSLRC Temperature Servo Turned OFF temporarily

Quote:

 

 My attempt to passively measure the transfer function of the foam failed fantastically.

As it turns out, the room temperature fluctuations inside the PSL box reach the 1 mK/rHz noise floor of the  AD590 (or maybe the ADC) at ~1-2 mHz. Everything at higher frequencies is noise.

So to see what the foam is doing we will have to do something smarter - we need a volunteer to disable the RC temperature servo from the EPICS screen and then cycle the PSL table lights every hour in the morning.

We'll then use our knowledge of the Laplace transform to get the TF from the step responses.

 more detailed instructions needed....

  2875   Tue May 4 02:28:38 2010 ranaConfigurationPSLRC Temperature Servo Turned OFF temporarily

 

 My attempt to passively measure the transfer function of the foam failed fantastically.

As it turns out, the room temperature fluctuations inside the PSL box reach the 1 mK/rHz noise floor of the  AD590 (or maybe the ADC) at ~1-2 mHz. Everything at higher frequencies is noise.

So to see what the foam is doing we will have to do something smarter - we need a volunteer to disable the RC temperature servo from the EPICS screen and then cycle the PSL table lights every hour in the morning.

We'll then use our knowledge of the Laplace transform to get the TF from the step responses.

  2874   Mon May 3 19:21:43 2010 AlbertoDAQEnvironmentBoot fest

[Alberto, Koji, Rana]

The RFM network failed today. We had to reboot the frame builder anf restart all the front end following the instructions for the "Nuclear Option".

Burt-restoring to May 1st at 18:07, or April 30 18:07 made c1sosvme crash. We had to reset the front ends again and restore to April 15th at 18:07 in order to make everything work.

Everything seems fine again now.

  2873   Mon May 3 17:49:41 2010 ranaConfigurationPSLRC Temperature Servo Turned OFF temporarily

Quote:

In order to measure the transfer function of the RC cavity's foam, I've turned off the servo so that the room temperature noise can excite it.

The attached plot shows a step response test from 2 weeks ago. Servo is nominally still working fine.

 I've just now re-enabled the temperature control of the reference cavity can. Trend of the last 8 days is attached.

Attachment 1: rct.png
rct.png
  2872   Mon May 3 16:53:27 2010 josephbUpdateCDSUpdated lsc.mdl and the ifo plant model with memory locations

I've updated the LSC and IFO models that Rana created with new shared memory locations.  I've used the C1:IFO- for the ifo.mdl file outputs, which in turn are read by the lsc.mdl file.  The LSC outputs being lsc control signals are using C1:LSC-.  Optics positions would presumably be coming from the associated suspension model, and am currently using SUP, SPX, and SPY for the suspension plant models (suspension vertex, suspension x end, suspension y end).

I've updated the web view of these models on nodus.  They can be viewed at: https://nodus.ligo.caltech.edu:30889/FE/

I've also created a C1.ipc file in /cvs/cds/caltech/chans/ipc  which assigns ipcNum to each of these new channels in shared memory.

  2871   Mon May 3 15:39:39 2010 josephbUpdateCDSDaily Downs update

Talked with Jay briefly today.  Apparently there are 3 IO chassis currently on the test stand at Downs and undergoing testing (or at least they were when Alex and Rolf were around).  They are being tested to determine which slots refer to which ADC, among other things. Apparently the numbering scheme isn't as simple as 0 on the left, and going 1,2,3,4, etc.  As Rolf and Alex are away this week, it is unlikely we'll get them before their return date.

Two other chassis (which apparently is one more than the last time I talked with Jay), are still missing cards for communicating between the computer and the IO chassis, although Gary thinks I may have taken them with me in a box.  I've done a look of all the CDS stuff I know of here at the 40m and have not seen the cards.  I'll be checking in with him tomorrow to figure out when (and if) I have the the cards needed.

  2870   Mon May 3 01:35:41 2010 KojiUpdateSUSLessons learned from MC spot centering

Lessons learned on the beam spot centering (so far)

Well-known fact:

The spot position on MC2 can be adjusted by the alignment of the mirror while maintaining the best overlapping between the beam and the cavity axes.

In general, there are two methods:

1) Use the cavity as a reference:
Move the MC mirrors such that the cavity eigenmode hits the centers of the mirrors.
-> Then adjust the incident beam to obtain the best overlapping to the cavity.

2) Use the beam as a reference:
Move the incident beam such that the aligned cavity has the spots at the centers of the mirrors.
-> Then adjust the incident beam to obtain the best spot position while the cavity mirrors keep tracking
the incident beam.

Found the method 1) is not practical.

This is because we can move the eigenmode of the cavity only by very tiny amount if we try to keep the cavity locked.
How much we can move by mirror alignment is smaller than the waist radius or the divergence angle.
For the MC, the waist radius is ~2mm, the divergence angle is 0.2mrad. This means the axis
translation of ~1mm is OK, but the axis rotation of ~4mrad is impractical.

Also it turned out that adjustinig steering mirror to the 10-m class cavity is quite difficult.
A single (minimum) touch of the steering mirror knob is 0.1mrad. This already change the beam position ~0.1mm.
This is not an enough resolution.

Method 2) is also not so easy: Steering mirrors have singular matrix

Indeed! (Remember the discussion for the IMMT)

What we need is the pure angle change of 4mrad at the waist which is ~2m distant from the steering mirror.
This means that the spot at the steering mirror must be moved by 8mm (= 4mrad x 2m). This is the result of the
nearly-singular matrix of the steering mirrors.

We try to avoid this problem by moving the in-vac mirror (IM1), which has somewhat independent move.
The refl beam path also has the big beam shift.
But once the vacuum manifold is evacuated we can adjust very little angle.

This can also be a good news: once the angle is set, we hardly can change it at the PSL side.

  2869   Mon May 3 01:16:50 2010 ranaHowToElectronicsMarconi phase noise measurement setup

 To try the 3-corner hat method on the Marconis, I started to set up the measurement into the DAQ system.

I have set the bottom 2 in the PSL rack to 11.1 MHz. I use a ZP-3MH level 13 mixer as the phase detector. The top one is the LO, it has an output of +13 dBm.

The bottom one is the test unit, it has an output of +6 dBm (should be close to the right level - the IP3 point is +9 dBm). The top one has external DC FM modulation enabled with a FM dev range of 10 Hz.

Mixer output goes through a 50 Ohm in-line termination and then a BLP-5 low pass filter (Steve, please order ~7 of the BLP-1.5 or BLP-1.9 low pass filter from Mini-Circuits) and then into

the DC coupled of a SR560. After some gain and filtering that feedback goes back to the FM input of the top-Marconi to close the PLL. I adjusted the gain to be as small as possible and still stay locked and not

saturate the ADC.

The input to the SR560 is Tee'd into another SR560 with AC coupled input, G = 1000, low-noise. Its output is going directly to the ADC channel - C1:IOO-MC_DRUM1.

I calibrated the channel by opening the loop and setting the AC coupled gain to 1. This lets the Marconis beat at several Hz. The peak-peak signal is equivalent to pi radians.

 

As usual, I was befuddled by the FM input. For some reason I always forget that since its a straight FM input, we don't need any filtering to get a plain 1/f loop. The attached plot shows how we get bad gain peaking if you forget this and use a 0.03 Hz pole in the SR560.

The grey trace is the ADC signal with everything hooked up, but the RF input set to zero (via setting Carrier = OFF in the bottom Marconi). It is the measurement noise.

The BLUE trace is very close to the true phase noise beat of the two Marconis with a calibration error of ~5%. I have not corrected for the loop gain: its right now around a 1 Hz UGF and 1/f. Next, I will measure the loop and compensate for it in the DTT calibration.

Then I'll measure the relative phase noise of 3 of the signal generators to get the individual noises.

Bottom line is that the sensitivity of this approach is good and we should do this rather that use spectrum analyzers since its easy to get very long averages and high res spectra. To get 5x better sensitivity, we can just use the Rai-FET box instead of a SR560 for the readout, but just have to contend with its batteries. Also should try using BALUNs on the RF and LO signals to get rid of the ground loops.

Attachment 1: pn.png
pn.png
  2868   Mon May 3 00:36:49 2010 KojiUpdateSUSHow to steer the incident beam to the MC?

Actually, I tried some tweaks of the input steering to get the beam being more centered on the MC mirrors on Saturday evening.

I made a mistake in the direction of the IM1Y tuning, and it made the horizontal spot position worse.

But, this also means that the opposite direction will certainly improve the horizontal beam angle.

Rotate IM1Y CCW!!!


The current setting is listed below

Alignment
MC1P 3.2531
MC1Y -0.5327
MC2P 3.3778
MC2Y -1.366
MC3P -0.5534
MC3Y -2.607


Spot positions
MC1H = +1.15 mm
MC1V = -0.13 mm
MC3H = +0.80 mm
MC3V = -0.20 mm

 

Quote:

Deviations of the MC spot from the center of the mirrors were measured.

MC1H = +0.29 mm
MC1V = -0.43 mm
MC3H = +1.16 mm
MC3V = -0.68 mm

 

  2867   Sun May 2 17:16:43 2010 KojiUpdateSUSHow to steer the incident beam to the MC?

Deviations of the MC spot from the center of the mirrors were measured.

MC1H = +0.29 mm
MC1V = -0.43 mm
MC3H = +1.16 mm
MC3V = -0.68 mm

1) The vertical deviation looks easy being adjusted as they are mostly translation. They are ~0.5mm too high.
The distance from SM2 to MC is 1.8m. Thus what we have to do is
rotate SM2 Pitch in CW knob by 0.25mrad.
1 turn steers the beam in 10mrad. So 0.25mrad is 1/40 turn (9deg)

2) The horizontal deviation is more troublesome. The common component is easily being adjusted
but the differential component (i.e. axis rotation) involves large displacement of the beam
at the periscope sterring mirrors.

(MC3H - MC1H) / 0.2 m * 1.8 m = 8 mm

The beam must be moved in 8mm at the periscope. This is too big.

We need to move the in-vac steering mirror IM1. Move SM2Yaw in 7mrad. This moves the spot on IM1 by 5mm*Sqrt(2).
Then Move Im1 Yaw such that we see the resonance.

For the alignment adjustment, try to maximize the transmission by MC2 Yaw (cavity axis rotation) and SM2Y (beam axis translation)  

Actual move will be:

- Move IM1Y CCW (assuming 100TPI 1.5 turn in total...half turn at once)
- Compensate the misalignment by SM2Y CW as far as possible.
- Take alignment with MC2Y and SM2Y as far as possible

This operation will move the end spot something like 15mm. This should be compensated by the alignment of MC1Y at some point.

Attachment 1: steering.png
steering.png
  2866   Sun May 2 16:52:44 2010 KojiSummarySUSCoil Actuator Balancing and Spot Position

Yes, of course. But so far I am trusting that the coils are inheretly balanced.
Probably you are talking about the dependence of the nodal position on the frequency...I need to check if 18Hz is sufficiently high or not for 0.1mm precision.

Also I am practicing myself to understand how I can adjust them by which screws as we probably have to do this adjustement many times.
(i.e. removal of the MZ, move of the table, PSL renewal and so on)

For the actuator calibration, we may be able to calibrate actuator responses by shaking them one by one while reading the OPLEV P/Y signals.

 

Quote:

 Oh, but it gets even better: in order to trust the A2L script in this regard you have to know that the coil driver - coil - magnet gain is the same for each channel. Which you can't.

But we have these handy f2pRatio scripts that Vuk and Dan Busby worked on. They use the optical levers to balance the actuators at high frequency so that the A2L gives you a true spot readout.

But wait! We have 4 coils and the optical lever only gives us 2 signal readouts...

 

  2865   Sun May 2 15:38:12 2010 ranaSummarySUSCoil Actuator Balancing and Spot Position

 

Oh, but it gets even better: in order to trust the A2L script in this regard you have to know that the coil driver - coil - magnet gain is the same for each channel. Which you can't.

But we have these handy f2pRatio scripts that Vuk and Dan Busby worked on. They use the optical levers to balance the actuators at high frequency so that the A2L gives you a true spot readout.

But wait! We have 4 coils and the optical lever only gives us 2 signal readouts...

  2864   Sun May 2 15:28:25 2010 KojiUpdateIOOSpot Positions of MC1/MC3

Summary

The spot positions on the MC mirrors were measured with coil balance gains.
The estimated spot positions from the center of the MC1 and MC3 are as followings:

MC1H = +0.29 mm
MC1V = -0.43 mm
MC3H = +1.16 mm
MC3V = -0.68 mm

The cordinates are described in the figure

Method

As far as the cavity mirrors are aligned to the incident beam, spots on the MC1 and MC3 tell us the geometry of the incident beam.
Note that spot position on the MC2 is determined by the alignment of the MC1 and MC3, so it does not a big issue now.
The calibration between the coil balance and the spot position are described in the previous entry.

  1. Lock the MC. Align it with MC2/MC3
  2. Run A2L scripts. script/A2L/A2L_MC1 and so on.
    • The scripts run only on the solaris machines. They require "expect" in stalled some specific place which does not exist on the linux machines.
    • Excitation amplitude, excitation freq, readback channels were modified

Result

Beam powers
MC Trans: 0.18
MC Refl: 0.12-0.13

Alignment biases
MC1P 3.2531
MC1Y -1.0827
MC2P 3.4534
MC2Y -1.1747
MC3P -0.9054
MC3Y -3.1393

Coil balances
MC1H 1.02682
MC1V 0.959605
MC3H 0.936519
MC3V 1.10755

(subtract 1, then multiply 10.8mm => spot position.)

Attachment 1: spot_position.png
spot_position.png
  2863   Sun May 2 13:04:51 2010 KojiSummarySUSCoil Actuator Balancing and Spot Position

I liked to know quantitatively where the spot is on a mirror.

With an interferometer and A2L scripts, one can make the balance of the coil actuators
so that the angle actuation does not couple to the longitudinal motion.
i.e. node of the rotation is on the spot

Suppose you have actuator balancing (1+α) f and (1-α) f.

=> d = 0.016 x α [m]

Full Imbalance   α = 1      -> d = 15 [mm]
10% Imbalance α = 0.1   -> d = 1.5 [mm]
1% Imbalance   α = 0.01 -> d = 0.15 [mm]


Eq of Motion:

I ω2 θ =  2 R f 
(correction) - I ω2 θ =  D f cos(arctan(L/2/D))
(re-correction on Sep 26, 2017) - I ω2 θ =  D f

m ω2 x = 2 α f ,
(correction) - m ω2 x = 2 α f ,

where R is the radius of the mirror, and D is the distance of the magnets. (kinda D=sqrt(2) R)

d, position of the node distant from the center, is given by

d = x/θ = α I / (m R) = 2 α β / D,

where β is the ratio of I and m. Putting R=37.5 [mm], L=25 [mm], β = 4.04 10-4 [m2], D~R Sqrt(2)

i.e. d = 0.015 α [m]

Attachment 1: coil_balance.png
coil_balance.png
  2862   Fri Apr 30 23:16:51 2010 KojiUpdateSUSSRM/PRM ready for baking

Kiwamu and Koji

- Checked the SRM/PRM balancing after the gluing.

- The mirrors were removed from the suspensions for baking.

- Bob is going to bake them next week.

  2861   Thu Apr 29 15:48:47 2010 josephbUpdateCDSNew CDS overview diagram in wiki

I've added a diagram in the wiki under IFO Upgrade 2009-2010->New CDS->Diagram section Joe_CDS_Plan.pdf (the .svg file I used to create it is also there).  This was mostly an exercise in me learning inkscape as well as putting out a diagram with which lists control and model names and where they're running.

A direct link is: CDS_Plan.pdf

  2860   Thu Apr 29 14:37:16 2010 josephbUpdateCDSNew Channel Name to Memory Location file

Awhile back we had requested a feature for the RCG code where a single file would define a memory location's name as well as its explicit hex address.  Alex told me it had been implemented in the latest code in SVN.  After being unable to find said file, I went back and talked to him and Rolf.  Rolf said it existed, but had not been checked into the SVN yet. 

I now have a copy of that file, called G1.ipc.  It is supposed to live in /cvs/cds/caltech/chans/ipc/ , so I created the ipc directory there.  The G1.ipc file is actually for a geo install, so we'll eventually make a C1.ipc file.

The first couple lines look like:

# /cvs/cds/geo/chans/ipc/G1.ipc
[default]
ipcType=SHMEM
ipcRate=2048
ipcNum=0
desc=default entry

[G1:OMC-QPD1P]
ipcType=SHMEM
ipcRate=32768
ipcNum=0
desc=Replaces 0x2000
#[G1:OMC-NOTUSED]
#ipcType=SHMEM
#ipcRate=32768
#ipcNum=1

[G1:OMC-QPD2P]
ipcType=SHMEM
ipcRate=32768
ipcNum=1
desc=Replaces 0x2008

 

There are also section using ipcType IPC:

[G1:SUS-ADC_CH_24]
ipcType=PCI
ipcRate=16384
ipcNum=1
desc=Replaces 0x20F0
[G1:SUS-ADC_CH_25]
ipcType=PCI
ipcRate=16384
ipcNum=2
desc=Replaces 0x20F0

 

Effectively the ipcNum tells it which memory location to use, starting with 0x2000 (at least thats how I'm interpreting it.  Every entry of a given ipcType has a different ipcNum which seems to be correlated to its description (at least early on - later in the file many desc= lines repeat, which I think means people were copy/pasting and got tired of editing the file.  Once I get a C1.ipc file going, it should make our .mdl files much more understandable, at least for communicating between models.  It also looks like it somehow interacts with the ADCs/DACs with ipcType PCI, although I'm hoping to get a full intro how to use the file tomorrow from Rolf and Alex.

  2859   Wed Apr 28 16:15:02 2010 KevinUpdatePSLAccelerometer Calibration

Koji, Steve, and Kevin looked into calibrating the Wilcoxon accelerometers. Once calibrated, the accelerometers will be used to monitor the motion of the PSL table.

We want to use the shaker to shake each accelerometer and monitor the motion with an OSEM. We will make a plate to attach an accelerometer to the shaker. A flag will also be mounted on this plate.The OSEM will be mounted on the table next to the shaker and positioned so that the flag can block the LED light as the plate moves up and down. We will then measure the motion of the accelerometer as it is shaken from the OSEM signal. The OSEM signal will be calibrated by keeping the plate and the flag still and moving the OSEM down along the flag a known distance with a micrometer.

  2858   Wed Apr 28 14:42:55 2010 ZachUpdateIOOMC alignment

Sure. I figured I would put up a How-To if it works. 

Quote:

 

 That's interesting.

Would it be possible to write about the technique on a wiki page as you get measurements and results?

 

  2857   Wed Apr 28 14:22:36 2010 KevinUpdatePSLre: 2W Vertical Beam Profile

I used the Mathematica CurveFit package that we use in Ph6/7 to make the fits for the beam profile data. I wrote two functions that use CurveFit shown in the attachment to make the fits to the error function and square root.

Attachment 1: BeamFit.nb.tar
  2856   Wed Apr 28 14:15:58 2010 AlbertoUpdateIOOMC alignment

 

 That's interesting.

Would it be possible to write about the technique on a wiki page as you get measurements and results?

  2855   Wed Apr 28 12:05:44 2010 ZachUpdateIOOMC alignment

I have worked out the first set of adjustments to make on the MC mirrors (all angle figures are in units of the increments on the control screen)

Using the method described in the previous post, I obtained the following matrix relating the angle-to-length coupling and the angular deviations. In the following matrix, Mij corresponds to the contribution of the jth degree of freedom to the ith A-to-L coupling, with the state vector defined as xi = (MC1P, MC2P, MC3P, MC1Y, MC2Y, MC3Y), where each element is understood as the angular deviation of the specific mirror in the specific direction from the ideal position, such that x = 0 when the cavity eigenmode is the correct one and the beams are centered on the mirrors (thus giving no A-to-L coupling regardless of the components of M).

 

M =

   1.0e+03 *

   -0.2843   -0.4279   -0.1254         0         0         0

   -0.8903   -0.4820   -0.6623         0         0         0

    0.5024    0.0484   -0.0099         0         0         0

         0         0         0         0.1145   -0.1941   -0.3407

         0         0         0         0.0265    1.5601    0.2115

         0         0         0         0.1015    0.1805   -0.0103,

giving an inverse

M-1 =  

    0.0003   -0.0001    0.0020         0         0         0

   -0.0031    0.0006   -0.0007         0         0         0

    0.0018   -0.0018   -0.0022         0         0         0

         0         0         0        -0.0013   -0.0015    0.0117

         0         0         0         0.0005    0.0008   -0.0008

         0         0         0        -0.0037   -0.0010    0.0044

The initial coupling vector is then acted on with this inverse matrix to give an approximate state vector x containing the angular misalignments of each mirror in pitch and yaw. The results are below:

x

   1P:  0.0223

   2P: -0.0733

   3P:  0.3010

  1Y:  -0.1372

  2Y:   0.0194

  3Y:  -0.0681

 

  2854   Wed Apr 28 09:21:16 2010 ZachUpdateelogelog restarted

And again.

Quote:

 Restarted the elog with the script as it was down.

 

  2853   Wed Apr 28 08:55:19 2010 ZachUpdateelogelog restarted

 Restarted the elog with the script as it was down.

  2852   Tue Apr 27 22:28:58 2010 ZachUpdateIOOMC alignment

Beginning last week, I have been helping Koji with some of the IO work that must be done for the 40m upgrade. The first thing he asked me to do is to help with the alignment of the MC.

As I understand, it became apparent that the IFO beam was not centered on all (or any) of the MC mirrors, which is disadvantageous for obvious reasons. We are trying to correct this, using the following strategy:

  1. Adjust the MC mirrors into rough alignment, isolate a strong TEM00, and lock the cavity
  2. Fine-tune the alignment by minimizing the REFL power when locked (in these first two steps, we adjusted only MC2 & MC3, assuming that the REFL beam was centered on the PD, and wanting to keep it that way). At this point, the cavity is resonating some asymmetric mode, looking something like (not to scale---for illustration only):MC_misaligned.png
  3. Shake all three mirrors (in succession) in pitch and yaw, each time demodulating the error signal at the frequency used for the excitation and recording the magnitude and phase of the response.
  4. Move one mirror's DC orientation, repeat step 3, and then restore the mirror to its original position
  5. Repeat step 4 for both angular degrees of freedom of each mirror

Using the results of these measurements, it is possible to evaluate the components of a block-diagonal matrix M which relates the tilt-to-displacement coupling of each DOF to each mirror's misalignment in that degree, i.e.,

a = M x

with a a 6-dimensional vector containing the coupling of each degree of freedom to the length of the cavity and x a 6-dimensional vector containing the angular misalignments of each. Due to orthogonality of pitch and yaw, M will take the form of a 6x6 matrix with two non-zero 3x3 blocks along the diagonal and zero matrices on the off-diagonal blocks.

The idea is to isolate components of M by moving one mirror at a time, solve for them, then find the inverse M-1 that should give us the required angular adjustments to obtain the beam-centered ideal cavity mode.

In theory, this need only be done once; in practice, our measurement error will compound and M will not be accurate enough to get the beams exactly centered, so we will have to iterate.

NOTE: The fact that we are adjusting the three cavity mirrors to obtain the ideal mode means that we will necessarily tarnish our coupling into the cavity. Once we have adjusted the mirrors once, we will need to re-steer the input beam and center it on the REFL diode. 

Status: This process has been completed once through step 5. I am in the process of trying to construct the matrix for the first adjustment.

 

  2851   Tue Apr 27 15:29:16 2010 KevinUpdatePSLre: 2W Vertical Beam Profile

I thought that the micrometer I was using to move the razor through the laser beam was metric; however, it is actually english.

After discovering this mistake, I converted my previous measurements to centimeters and fit the data to

w = sqrt(w0^2+lambda^2*(z-z0)^2/(pi*w0)^2) with the following results:

reduced chi squared = 14.94

z0 = (-4.2 ± 1.9) cm

w0 = (0.013 ± 0.001) cm

Attachment 1: vbp.jpg
vbp.jpg
Attachment 2: vbp_residuals.jpg
vbp_residuals.jpg
  2850   Tue Apr 27 14:18:53 2010 kiwamuUpdateGreen Lockingwaist positon of Gaussian beam in PPKTP crystals

The mode profile of Gaussian beams in our PPKTP crystals was calculated.

I confirmed that the Rayleigh range of the incoming beam (1064 nm) and that of the outgoing beam (532 nm) is the same.

And it turned out that the waist postion for the incoming beam and the outgoing beam should be different by 13.4 mm toward the direction of propagation.

These facts will help us making optical layouts precisely for our green locking.


(detail)

The result is shown in the attached figure, which is essentially the same as the previous one (see the entry).

The horizontal axis is the length of the propagation direction, the vertical axis is the waist size of Gaussian beams.

Here I put x=0 as the entering surface of the crystal, and x=30 mm as the other surface.

The red and green solid curve represent the incoming beam and the outgoing beam respectively. They are supposed to  propagate in free space.

And the dashed curve represents the beams inside the crystal.

A trick in this calculation is that: we can assume that  the waist size of 532 nm is equal to that of 1064 nm divided by sqrt(2) . 

If you want to know about this treatment in detail,  you can find some descriptions in this paper;

"Third-harmonic generation by use of focused Gaussian beams in an optical super lattice" J.Opt.Soc.Am.B 20,360 (2003)"

Attachment 1: mode_in_PPKTP.png
mode_in_PPKTP.png
  2849   Tue Apr 27 11:16:13 2010 josephbConfigurationCDSWiki page with CDS .mdl names, shared memory allocation

I've added a new page in the wiki which describes the current naming scheme for the .mdl model files used for the real time code generator.  Note, that these model names do not necessarily have to be the names of the channels contained within.  Its still possible to make all suspension related channels start with C1:SUS- for example.  I'm also allocating 1024 8 byte channels for shared memory address space for each controller and each simulated plant.

The wiki page is here

Name suggestions, other front end models that are needed long term (HEPI is listed for example, even though we don't have it here, since in the long run we'd like to port the simulated plant work to the sites) are all welcome.

  2848   Mon Apr 26 21:12:53 2010 KojiUpdateSUSPRM/SRM standoffs glued

Kiwamu and Koji

The PRM/SRM were balanced with the standoffs. We glued them to the mirror.

This was the last gluing so far until we get new PRM/ETMs.

  2847   Mon Apr 26 17:34:31 2010 KojiUpdatePSLre: 2W Vertical Beam Profile

Give me the plot of the fit, otherwise I am not convinced.

Quote:

I tried Koji's suggestions for improving the fit to the vertical beam profile; however, I could not improve the uncertainties in the fit parameters.

  2846   Mon Apr 26 16:51:37 2010 KevinUpdatePSLre: 2W Vertical Beam Profile

I tried Koji's suggestions for improving the fit to the vertical beam profile; however, I could not improve the uncertainties in the fit parameters.

I started retaking the data today with the same laser settings used last time and noticed that the photodiode was saturating. We were using an ND 4.0 neutral density filter on the photodiode. Koji and I noticed that the coating on the filter was reduced in the center and added an additional ND 0.6 filter to the photodiode. This seemed to fix the photodiode saturation.

I think that the photodiode was also saturating to a lesser extent when I took the last set of data. I will take another vertical beam profile tomorrow.

[Edit by KA: Metallic coating started being evaporated and the ND filters reduced their attenuation. We decided to use absorptive one as the first incident filter, and put a thinner one behind. This looked fine.]

  2845   Mon Apr 26 12:24:58 2010 josephbUpdateGeneralDaily Downs update

Talked with Jay briefly this morning.

We are due another 1-U 4 core (8 CPU) machine, which is one of the ones currently in the test stand.  I'm hoping sometime this week I can convince Alex to help me remove it from said test stand.

The megatron machine we have is definitely going to be used in the 40m upgrade (to answer a question of Rana's from last Wednesday's meeting).  Thats apparently the only machine of that class we get, so moving it to the vertex for use as the LSC or SUS vertex machine may make sense.  Overall we'll have the ASS, OMC, Megatron (SUS?), along with the new 4 1-U machines, for LSC, IO, End Y and End X.  We are getting 4 more IO chassis, for a total 5.  ASS and OMC machine will be going without full new chassis.

Speaking of IO chassis, they are still being worked on.  Still need a few cards put in and some wiring work done.  I also didn't see any other adapter boards finished either.

  2844   Mon Apr 26 11:29:37 2010 josephbUpdateComputersUpdated bitwise.pm in RCG SVN plus other fixes

To fix a problem one of the models was having, I checked the CVS version of the Bitwise.pm file into the SVN (located in /home/controls/cds/advLigoRTS/src/epics/util/lib), which adds left and right bit shifting funtionality.  The yec model now builds with the SVN checkout.

Also while trying to get things to work, I discovered the cdsRfmIO piece (used to read and write to the RFM card) now only accepts 8 bit offsets.  This means we're going to have to change virtually all of the RFM memory locations for the various channels, rather than using the values from its previous incarnation, since most were 4 bit numbers.  It also means it going to eat up roughly twice as much space, as far as I can tell.

Turns out the problem we were having getting to compile was nicely answered by Koji's elog post.  The shmem_daq value was not set to be equal to 1.  This caused it to look for myrimnet header files which did not exist, and caused compile time errors.  The model now compiles on megatron.

[Edit by KA: 4 bit and 8 bit would mean "bytes". I don't recall which e-log of mine Joe is referring.]

 

  2843   Mon Apr 26 11:14:04 2010 KojiUpdateGreen LockingTemperature scan for PPKTP

I scanned the temperature of the crystal oven on Friday night in order that we can find the optimal temperature of the crystal for SHG.

The optimal temperature for this crystal was found to be 36.2 deg.


The crystal is on the PSL table. The incident beam on the crystal is 27.0mW with the Newport power-meter configured for 1064nm.
The outgoing beam had 26.5mW.

The outgoing beam was filtered by Y1-45S to eliminate 1064nm. According to Mott's measurements, Y1-45S has 0.5% transmission for 1064nm, while 90% transmission for 532nm. This means I still had ~100uW after the Y1-45S. This is somewhat consistent with the offset seen in the power-meter reading.

First, I scanned the temperature from 28deg to 40deg with 1deg interval.The temperature was scaned by changing the set point on the temperature controller TC-200.The measurements were done with the temperature were running. So, the crystal may have been thermally non-equilibrium.

Later, I cut the heater output so that the temperature could be falling down slowly for the finer scan. The measurement was done from 38deg to 34deg with interval of 0.1deg with the temperature running.

I clearly see the brightness of the green increase at around 36 deg. The data also shows the peak centered at 36.2deg. We also find two lobes at 30deg and 42deg. I am not sure how significant they are.

Attachment 1: SHG_pow.png
SHG_pow.png
  2842   Mon Apr 26 11:13:28 2010 steveOmnistructurePEMAC on/off is working

Quote:

Koji and I wanted to turn off the IFO-room AC so the wind would not blow on MC1-3. We could not. The switches were probably bypassed when the power transformer was replaced at the last scheduled power outage.

 There is one three position manual/off/auto switch next to the filter for each unit at CES. They have to be in AUTO position when we want to turn AC on/off from the lab.

  2841   Mon Apr 26 10:21:45 2010 josephbUpdateLSCStarted dev of LSC FE

Quote:

Joe and I started working on the new LSC FE control today. We made a diagram of the system in Simulink, but were unable to compile it.

Joe checked out the latest CDS software out of their new SVN and put it somewhere (perhaps his home directory).

The SVN checkout was done on megatron.  It is located under /home/controls/cds/advLigoRTS

So, to compile (or at least try to) you need to copy the .mdl file from /cvs/cds/caltech/cds/advLigo/src/epics/simLink to /home/controls/cds/advLigoRTS/src/epics/simLink on megatron, then run make SYS in the advLigoRTS directory on megatron.

The old checkout from CVS exists on megatron under /home/controls/cds/advLigo.

  2840   Sun Apr 25 10:40:21 2010 KojiUpdateLSCStarted dev of LSC FE

Once you made a CDS model, please update the following wiki page. This will eventually help you.

http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/Existing_RCG_DCUID_and_gds_ids

Quote:

LSC Plant Model. That is all.

 

  2839   Sun Apr 25 02:56:07 2010 ranaUpdateLSCStarted dev of LSC FE

LSC Plant Model. That is all.

  2838   Sat Apr 24 15:50:47 2010 KojiUpdatePSLre: 2W Vertical Beam Profile

1. The vertical axis should start from zero. The horizontal axis should be extended so that it includes the waist. See Zach's plot http://nodus.ligo.caltech.edu:8080/40m/2818

2. Even if you are measuring only the linear region, you can guess w0 and z0, in principle. w0 is determined by the divergence angle (pi w0/lambda) and z0 is determined by the linear profile and w0. Indeed your data have some fluctuation from the linear line. That could cause the fitting prescision to be worse.

3. Probably the biggest reason of the bad fitting would be that you are fitting with three parameters (w0, z0, zR) instead of two (w0, z0). Use the relation ship zR= pi w0^2/lambda.

  2837   Sat Apr 24 15:05:41 2010 KevinUpdatePSL2W Vertical Beam Profile

The vertical beam profile of the Innolight 2W laser was measured at eight points along the axis of the laser.

These measurements were made with the laser crystal temperature at 25.04°C and the injection current at 2.091A. z is the distance from the razor blade to the flat black face of the front of the laser.

The voltage from a photodiode was measured for the razor at a number of heights. Except for the first two points, one scan was made with the razor moving down and a second scan was made with the razor moving up. This data was fit to

y = a*erf(sqrt(2)*(x-x0)/w) + b with the following results:

z(cm) (±0.1cm) w(mm) chi^2/ndf
3.9 0.085 ± 0.006 77.09
6.4 0.130 ± 0.004 12.93
8.8 down 0.145 ± 0.008 66.57
8.8 up 0.147 ± 0.008 18.47
11.6 down 0.194 ± 0.010 64.16
11,6 up 0.214 ± 0.009 27.23
14.2 down 0.177 ± 0.008 49.95
14.2 up 0.183 ± 0.007 29.85
16.6 down 0.205 ± 0.006 18.35
16.2 up 0.203 ± 0.007 17.16
19.2 down 0.225 ± 0.007 18.92
19.2 up 0.238 ± 0.011 25.56
21.7 down 0.292 ± 0.006 11.30
21.7 up 0.307 ± 0.008 11.85

The values for w and its uncertainty were estimated with a weighted average between the two scans for the last six points and all eight points were fit to

w = w0*sqrt(1+(z-z0)2/zR2) with the following results:

chi^2/ndf = 17.88

w0 = (0.07 ± 0.13) mm

z0 = (-27 ± 121) mm

zR = (65 ± 93) mm

It looks like all of the data points were made in the linear region so it is hard to estimate these parameters with reasonable uncertainty.

Attachment 1: vbp.jpg
vbp.jpg
  2836   Fri Apr 23 21:02:14 2010 rana, joeUpdateLSCStarted dev of LSC FE

Joe and I started working on the new LSC FE control today. We made a diagram of the system in Simulink, but were unable to compile it.

Joe checked out the latest CDS software out of their new SVN and put it somewhere (perhaps his home directory).

We then copied the directory with the .mdl files and the CDS parts library into our real Simulink Model Directory:

/cvs/cds/caltech/cds/advLigo/src/epics/simLink

Use this and not someplace in Alex or Rob's home directory !

Joe will put in more details on Monday once he figures out how to build the new stuff. Basically, we decided not to support multiple versions of the CDS real time code here. We'll just stay synced to the latest stable ~versions.

I exported the current version of the LSC FE into our public_html/FE/ area on nodus where we will put all of the self-documenting FE diagrams:

https://nodus.ligo.caltech.edu:30889/FE/lsc_slwebview_files/index.html

To make a web setup like this, you just use the "Export to Web" feature from the top-level Simulink diagram (e.g. lsc.mdl). Choose the following options:

Untitled.png

Note: in order to get the web page to work, I had to change the apache httpd.conf file to allow AddType file overriding. Here's the term cap of the diff:

nodus:etc>diff httpd.conf httpd.conf~
155c155
< ServerAdmin jenne@caltech.edu
---
> ServerAdmin aso@caltech.edu
225d224
<     AllowOverride FileInfo

  2835   Fri Apr 23 18:30:49 2010 Aidan, Jenne, KojiSummaryGreen LockingGreen means GO!

Jenne, Koji and I assembled the Covesion Oven today, inserted a PPKTP crystal from Raicol, aligned the crystal to a 50mW focus and
got some green beam coming out.

Covesion Oven assembly

The oven contains a brass clip that can clamp a crystal up to 10mm wide x 0.5mm high x 40mm long (according to the instructions). According to the correspondence from Covesion the clip can accomodate a crystal up to 1.5mm high. Our crystal is 1mm x 1mm x 30mm.

  1. We removed the brass springs from the clip - see Koji's photos
  2. We placed the Raicol PPKTP crystal (#725) into the clamp with the long polished surfaces facing out to the sides and the roughened surfaces facing up and down.
  3. We balanced the 10mm x 40mm x 1mm glass plate on top of the crystal.
  4. We replaced the brass springs in the top of the clip but only tightened the screws a couple of turns so they wouldn't fall out.
  5. Very carefully and slowly, I tightened the screws a few turns in a star-shaped order to distribute the pressure evenly across the glass top
  6. Each time I tightened all eight screws, I jiggled each of the four springs to see if there was any compression in them
  7. Once the springs started to show signs of compression I stopped tightening them and tested the stability of the glass plate - a reasonable amount of pressure was required to move the plate - about the same amount required to push a SR560 across an optical table with your index finger.
  8. We loosely attached the lid and moved the oven to the table

Alignment of the crystal to the focus

The oven was mounted on a 4-axis Newport translation stage. We plonked the assembly onto the table, removed the lid and adjusted the rough position so that a focus of the 1064nm beam, from a 100mm lens, was positioned near the center of the crystal - then it was clamped down to the table. From here we adjusted the alignment of the stage, using an IR card and a viewer to guide us, until we eventually saw some green beam coming out. We were all very excited by this! We optimized the alignment as best we could using the IR card and then we replaced the lid on the oven. At this point the temperature of the PPKTP was around 26.5C and the green beam coming out look quite dim. We turned the oven up to around 36 degC and observed the beam getting much brighter and we approached the optimum phase-matching condition.

We haven't done anyway quantitative measurements yet but we were pleased with how easy this first stage was.

 

[Edit by Koji] More photos are on Picasa album

Attachment 1: IMG_2405.jpg
IMG_2405.jpg
Attachment 2: IMG_2417.jpg
IMG_2417.jpg
  2834   Thu Apr 22 21:42:24 2010 AlbertoUpdatePSLInnolight 2W Vertical Beam Profile

 

 What kind of fit did you use? How are the uncertainties in the parameters obtained?

  2833   Thu Apr 22 20:28:40 2010 JenneUpdatePEMSeismometer Noise Characteristics

Quote from elog 2237, 11 Nov 2009:

The attached plot shows the spectra of the 3 Z axes of the 3 seismometers we have (this data is from ~20Aug2009, when the Ranger was in the Z orientation) in Magenta, Cyan and Green, and the noise of each of the sensors in Red, Blue and Black.  The noise curves were extracted from the spectra using the Huddle Test / 3 Corner Hat method.  The Blue and Black traces which are just a few points are estimates of the noise from other spectra.  The Blue points come from the Guralp Spec Sheet, and the Black comes from the noise test that Rana and I did the other day with the Ranger (elog 2223).  

I'm not really happy with the black spectra - it looks way too high.  I'm still investigating to see if this is a problem with my calibration/method....

 So, as it turns out (surprise), I'm a spaz and forgot a 2*pi when calibrating the Guralp noise spectra from the spec sheet.  I noticed this when redoing the Huddle Test, and comparing my Spec Sheet Guralp noise with Rana's, which he shows in elog 2689.  When going from m/s^2, the units in the spec sheet, I just tilted the line by a factor of frequency.  Koji pointed out that I needed a factor of 2*pi*f.  That moves the Guralp spec line in the plot in elog 2237 (to which this entry is a reply) down by ~6, so that my measured noise is not, in fact, below the spec.  This makes things much more right with the world.

In other news, I redid the Huddle analysis of the 2 Guralp seismometers, ala Rana's elog 2689. The difference is now we are on the granite slab, with soft rubber feet between the floor and the granite.  We have not yet cut holes in the linoleum (which we'll do so that we're sitting directly on the 40m's slab).

Rana> this seems horrible. Its like there's a monster in there at 6-7 Hz! Either the seismos are not centered or the rubber balls are bad or Steve is dancing on the granite slab again.

 

Attachment 1: Gur1_Gur2_noise.png
Gur1_Gur2_noise.png
  2832   Thu Apr 22 15:44:34 2010 josephbUpdateComputers 

I updated the default FILTER.adl file located in /home/controls/cds/advLigo/src/epics/util/ on megatron.  I moved the yellow ! button up slightly, and fixed the time string in the upper right.

Attachment 1: New_example_CDS_filter.png
New_example_CDS_filter.png
  2831   Thu Apr 22 09:03:54 2010 steveOmnistructurePEMaircondition can not be turned off

Koji and I wanted to turn off the IFO-room AC so the wind would not blow on MC1-3. We could not. The switches were probably bypassed when the power transformer was replaced at the last scheduled power outage.

ELOG V3.1.3-